Order the following from least to greatest. please show all of your work so i know how to do it thank you.
-3
-3/10
4/5
-2
0.5

Answer by stanbon(75887)

Order the following from least to greatest. please show all of your work so i know how to do it thank you.
--------------
Change each number to decimal form:
------
-3.0
-3/10 = -0.30
4/5 = 0.80
-2.0
0.50
----
Order: -3,-2,-0.3,0.5,0.8
So order:-3,-2,-3/10,0.5,4/5
===============================
Cheers,
Stan H.